The kinetic energy of a particle moving along a circle of radius R depends on the distance covered s as T = as2, where a is a constant. Find the force acting on the particle as a function of s.
1
Expert's answer
2020-10-19T13:19:22-0400

Explanations & Calculations


  • If the mentioned formula "\small T = as^2" shows the kinetic energy of the particle as "\small E_k = as^2" then by applying the Newton's second law towards the center of the particle's path the force acting on it could easily be calculated.
  • When a particle moves on a circular path, a particular force is needed in order to stick on to that path or it moves on a straight line & that force is called the centripetal force ("\small F_c").
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small F_c &= ma\\
&= \small m\frac{v^2}{R}\\
&= \small \frac{2}{R}\times \frac{1}{2}mv^2 : \text{rearrange to form the expression for} E_k\\
&= \small \frac{2}{R}\times as^2\\
F_c&= \small \bold{\frac{2as^2}{R}} 
\end{aligned}"


  • But note that it's said that the kinetic is not a constant as it is subjected to change as the covered distance changes meaning that there may be it's speed changes over time hence the velocity change both in radial & tangential directions.
  • Therefore, a tangential acceleration is present hence a tangential force therefore a resultant force to be calculated.
  • Tangential acceleration

"\qquad\qquad
\begin{aligned}
\small a_t &= \small \frac{dv}{dt}
\end{aligned}"

  • As its known from the given relationship

"\qquad\qquad
\begin{aligned}
\small \frac{1}{2}mv^2&= \small as^2\\
\small v^2&= \small \frac{2as^2}{m}\\
\small 2v\frac{dv}{dt}&= \small \frac{4as}{m}\frac{ds}{dt}=\frac{4as}{m}\times v\\
\therefore \small \frac{dv}{dt} = a_t&= \small \frac{2as}{m}
\end{aligned}"

  • Therefore, the tangential force "\small F_t = ma_t= \bold{2as}"
  • These forces are normal to each other hence the resultant is

"\qquad\qquad
\begin{aligned}
\small F_{r} &= \small \sqrt{F_r^2+F_t^2}\\
&= \small \sqrt{\frac{4a^2s^4}{R^2}+4a^2s^2}\\
F_r&= \small \bold{\frac{2as}{R}\sqrt{R^2+s^2}}
\end{aligned}"